### v1

# Note on the Humbert function $\Psi_{1}$

Abstract.
Recently, Hang and Luo (2024) used an elementary approach to derive the leading asymptotic behavior of the Humbert function $\Psi_{1}[a,b;c,c^{\prime};x,y]$ when $x\to\infty$ and $y\to+\infty$. In this paper, we give detailed asymptotic analyses of $\Psi_{1}$ when $x\to\infty$ and $y\to\infty$. Our proof is based on a uniform estimate of ${}_{2}F_{2}[a,b-n;c,d-n;z]$ for $z$ large.
Key words and phrases:
Humbert function, asymptotics, generalized hypergeometric function
2020 Mathematics Subject Classification:
33C20, 33C70, 41A60, 44A20
1. Introduction
The confluent hypergeometric function of two variables $\Psi_{1}$, introduced by Humbert [5, p. 75], is defined by
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\sum_{m,n=0}^{\infty}\frac{(a)_{m+n}(b)_{m}}{(c)_{m}(c^{\prime})_{n}}\frac{x^{m}}{m!}\frac{y^{n}}{n!},\qquad|x|<1,\quad|y|<\infty,$
where $a,b\in\mathbb{C}$ and $c,c^{\prime}\notin\mathbb{Z}_{\leqslant 0}$. It can be analytically extended to the region
$\mathbb{D}_{\Psi_{1}}:=\bigl{\{}(x,y)\in\mathbb{C}^{2}\colon x\neq 1,\,|{\arg}(1-x)|<\pi,\,|y|<\infty\bigr{\}}$
via the series representation ([5, p. 75]; see also [3, equation (1.2)])
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\sum_{n=0}^{\infty}\frac{\left(a\right)_{n}}{\left(c^{\prime}\right)_{n}}\,_{2}F_{1}\left[\begin{matrix}a+n,b\\ c\end{matrix};x\right]\frac{y^{n}}{n!},$
where ${}_{2}F_{1}$ denotes the Gauss hypergeometric function defined in (1.5).
By using inverse Laplace transformations of the Humbert functions $\Phi_{2}$, $\Phi_{3}$ and $\Xi_{2}$, Wald and Henkel [11] derived the leading-order behavior of these functions when two variables become simultaneously large. But they failed to render the integral for $\Psi_{1}$ [11, equation (2.4b)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{1}{\Gamma(a)}\int_{0}^{\infty}\mathrm{e}^{-u}u^{a-1}\,_{1}F_{1}\left[\begin{matrix}b\\ c\end{matrix};xu\right]{}_{0}F_{1}\left[\begin{matrix}-\\ c^{\prime}\end{matrix};yu\right]\mathrm{d}u$
as convolutions and so could not establish the asymptotics of $\Psi_{1}$.
Recently, Hang and Luo [3] partially answered Wald and Henkel’s problem by deriving the leading behavior of $\Psi_{1}$ under the condition
$x\to\infty,y\to+\infty,\quad\left|\arg(1-x)\right|<\pi,\quad 0<\gamma_{1}\leqslant\frac{y}{|1-x|}\leqslant\gamma_{2}<\infty.$ (1.1)
Their starting point is the following expression for $\Psi_{1}$.
Theorem 1.1 ([3, Theorem 3.1]).
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c\in\mathbb{C}\setminus\mathbb{Z}$. Then
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}\left(1-x\right)^{-a}V_{1}(x,y)+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)}\left(1-x\right)^{-b}V_{2}(x,y)$ (1.2)
holds for $|{\arg}(1-x)|<\pi$, $|x-1|>1$ and $|y|<\infty$, where
$V_{1}(x,y):=\sum_{n=0}^{\infty}\frac{\left(a\right)_{n}\left(c-b\right)_{n}}{\left(a-b+1\right)_{n}}{}_{2}F_{2}\left[\begin{matrix}a-c+1,a+n\\ c^{\prime},a-b+1+n\end{matrix};\frac{y}{1-x}\right]\frac{\left(1-x\right)^{-n}}{n!},$
$V_{2}(x,y):=\sum_{n=0}^{\infty}\frac{\left(b\right)_{n}\left(c-a\right)_{n}}{\left(b-a+1\right)_{n}}{}_{2}F_{2}\left[\begin{matrix}a-c+1,a-b-n\\ c^{\prime},a-c+1-n\end{matrix};y\right]\frac{\left(1-x\right)^{-n}}{n!}.$
In this paper, we shall give detailed asymptotic analyses of $\Psi_{1}$ when $x\to\infty$ and $y\to\infty$. Our first result provides the full asymptotic expansion of $\Psi_{1}$ under the condition (1.1), which covers Hang and Luo’s result [3, Theorem 3.6].
Theorem 1.2.
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c\in\mathbb{C}\setminus\mathbb{Z}$. Then, under the condition (1.1),
$\Psi_{1}[a,b;c,c^{\prime};x,y]\sim\frac{\Gamma(c)\Gamma(c^{\prime})}{\Gamma(a)\Gamma(c-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{\infty}a_{k}(x,y)y^{-k},$
where $a_{0}(x,y)=1$ and in general,
$a_{k}(x,y)=\sum_{j=0}^{k}\frac{\left(b\right)_{j}\left(b-a+c^{\prime}\right)_{k-j}\left(j+b-a+1\right)_{k-j}}{j!(k-j)!}{}_{3}F_{2}\left[\begin{matrix}-j,j-k,c+c^{\prime}-a-1\\ b-a+c^{\prime},a-b-k\end{matrix};1\right]\Big{(}\frac{y}{1-x}\Big{)}^{b+j}.$ (1.3)
Our second result gives the full asymptotics of $\Psi_{1}$ for two large arguments in the remaining cases. These two results give a complete answer to Wald and Henkel’s problem.
Theorem 1.3.
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c,b-c\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\{\Re(a-b),\Re(a-c)+1\}$ and that the fractional parts of $w-\Re(a-b)$ and $w-\Re(a-c)-1$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is small.
Then, under the condition
$x\to\infty,y\to\infty,\quad\left|\arg(1-x)\right|<\pi,\quad\left|\arg(-y)\right|<\pi,\quad 0<\gamma_{1}\leqslant\Big{|}\frac{y}{1-x}\Big{|}\leqslant\gamma_{2}<\infty$ (1.4)
and under the restriction that $y$ is bounded away from the points $b-a+k\,(k\in\mathbb{Z})$, the function $\Psi_{1}\equiv\Psi_{1}[a,b;c,c^{\prime};x,y]$ admits the asymptotic expansion
$\Psi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}(x,y)+\frac{\Gamma(c)\Gamma(c^{\prime})\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)\Gamma(b-a+c^{\prime})}(A_{2}(x,y)+A_{3}(x,y))$
$+\frac{\Gamma(c)\Gamma(c^{\prime})}{\Gamma(a)\Gamma(c-b)}A_{4}(x,y)+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-2b-c^{\prime})-N}\mathrm{e}^{\Re(y)}\right),$
where
$A_{1}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(a\right)_{k}\left(c-b\right)_{k}}{\left(a-b+1\right)_{k}k!}\,_{2}F_{2}\left[\begin{matrix}a-c+1,a+k\\ c^{\prime},a-b+1+k\end{matrix};\frac{y}{1-x}\right]\left(1-x\right)^{-a-k},$
$A_{2}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(c-b\right)_{k}}{k!}\left(-y\right)^{-a-k}\sum_{j=0}^{k-1}\frac{\left(b\right)_{j}\left(a-b\right)_{k-j}\left(a-b-c^{\prime}+1\right)_{k-j}}{j!\left(c-b\right)_{k-j}}\Big{(}\frac{y}{x-1}\Big{)}^{b+j},$
$A_{3}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(b\right)_{k}\left(c-b\right)_{k}}{k!^{2}}\left(-y\right)^{-a-k}{}_{2}F_{3}\left[\begin{matrix}b+k,b-c+1\\ 1+k,b-a+1,b-a+c^{\prime}\end{matrix};\frac{y}{1-x}\right]\Big{(}\frac{y}{x-1}\Big{)}^{b+k},$
$A_{4}(x,y)={}$ $y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k},$
with $M=\lfloor w+\Re(b-a)\rfloor$, $N$ being any positive integer and $a_{k}(x,y)$ given by (1.3).
The paper are organised as follows. In Section 2, we demonstrate some estimates of the generalized hypergeometric functions ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ when $z$ approaches to infinity in any way, which are valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$. In Section 3, we use the tools in Section 2 to prove our main results. The paper concludes with further remarks in Section 4.
Notation. The number $C$ generically denotes a positive constant independent of the parameter $n$ and the variable $z$. By $f(n,z)=\mathcal{O}(a_{n}g(z))\,(z\in\Omega)$, we mean that there exists a constant $K>0$ independent of $n$ and $z$ such that
$|f(n,z)|\leqslant K|a_{n}||g(z)|,\quad n\in\mathbb{Z}_{\geqslant 0},\,z\in\Omega.$
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by
${}_{p}F_{q}\left[\begin{matrix}a_{1},\cdots,a_{p}\\ b_{1},\cdots,b_{q}\end{matrix};z\right]\equiv{}_{p}F_{q}[a_{1},\cdots,a_{p};b_{1},\cdots,b_{q};z]:=\sum_{n=0}^{\infty}\frac{(a_{1})_{n}\cdots(a_{p})_{n}}{(b_{1})_{n}\cdots(b_{q})_{n}}\frac{z^{n}}{n!},$ (1.5)
where $a_{1},\cdots,a_{p}\in\mathbb{C}$ and $b_{1},\cdots,b_{q}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Empty products are taken to be $1$.
2. Auxiliary results
In this section, we prove some auxiliary results, including three lemmas about explicit bounds for the ratio of two gamma functions: (i) the first is benifitial to bound the tail series; (ii) the second generalizes the simple bound of gamma ratio [8, equation (5.6.8)]; (iii) the third guarantees the uniformity of explicit bounds for ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ obtained in the ensuing theorems.
Lemma 2.1 ([3, Lemma 2.1]).
If $a\in\mathbb{C}$ and $b\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$, then
$\left|\frac{\left(a\right)_{n}}{\left(b\right)_{n}}\right|\leqslant C\left(n+1\right)^{\Re(a-b)},\qquad n\in\mathbb{Z}_{\geqslant 0}.$ (2.1)
Lemma 2.2.
If $\Re(b-a)\geqslant 1,\,\Re(a)\geqslant 0$ and $\Re(z)>|\Im(a)|$, then
$\left|\frac{\Gamma(z+a)}{\Gamma(z+b)}\right|\leqslant\frac{\Gamma(\Re(b-a))}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\left(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta\right)^{\Re(a-b)},$
where $\theta=\arg(z)\in(-\frac{\pi}{2},\frac{\pi}{2})$.
Proof.
Assume that $\Re(b-a)\geqslant 1,\,\Re(a)\geqslant 0$ and $\Re(z)>|\Im(a)|$. Then the quotient of two gamma functions has the integral representation [9, p. 33]
$\frac{\Gamma(z+a)}{\Gamma(z+b)}=\frac{\mathrm{e}^{-\mathrm{i}\theta}}{\Gamma(b-a)}\int_{0}^{\infty}\mathrm{e}^{-|z|u-au\mathrm{e}^{-\mathrm{i}\theta}}\left(1-\mathrm{e}^{-u\mathrm{e}^{-\mathrm{i}\theta}}\right)^{b-a-1}\mathrm{d}u.$
Note that
$1-\mathrm{e}^{-u\mathrm{e}^{-\mathrm{i}\theta}}=1-\mathrm{e}^{-u\cos\theta}\cos(u\sin\theta)-\mathrm{i}\cdot\mathrm{e}^{-u\cos\theta}\sin(u\sin\theta)$
lies in the right half-plane and $|1-\mathrm{e}^{-\zeta}|\leqslant|\zeta|$ when $\left|\arg(\zeta)\right|\leqslant\frac{\pi}{2}$. Hence
$\left|\frac{\Gamma(z+a)}{\Gamma(z+b)}\right|$ $\leqslant\frac{1}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\int_{0}^{\infty}\mathrm{e}^{-(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta)u}u^{\Re(b-a)-1}\mathrm{d}u$
$=\frac{\Gamma(\Re(b-a))}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\left(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta\right)^{\Re(a-b)}.$
This completes the proof. ∎
Lemma 2.3.
If $a\in\mathbb{C}$ and $b\in\mathbb{C}\setminus\mathbb{Z}$, then
$\prod_{j=1}^{n}\left|\frac{-a+j-z}{-b+j-z}\right|\leqslant C\left(n+1\right)^{2|a-b|},\quad n\in\mathbb{Z}_{\geqslant 0}$
holds for $z$ bounded away from the points $-b+1,\cdots,-b+n$.
Proof.
Suppose that $n\geqslant 1$ and $z\in\bigcup_{j=1}^{n}\{|z+b-j|\geqslant\varepsilon\}$ for some $\varepsilon\in(0,\frac{1}{4})$. Denote
$P_{n}(z):=\prod_{j=1}^{n}\left|\frac{-a+j-z}{-b+j-z}\right|.$
Case 1. If $\Re(z+b)\in(k-\varepsilon,k+\varepsilon)$ for some $k\in\mathbb{Z}\cap[1,n]$, there is a point $z_{0}=-b+k+\varepsilon\,\mathrm{e}^{\mathrm{i}\beta}$ with $\Re(z)=\Re(z_{0})$, such that for $j\in\mathbb{Z}\cap[1,n]$,
$|-b+j-z|\geqslant|-b+j-z_{0}|.$
Therefore,
$P_{n}(z)$ $\leqslant\prod_{j=1}^{n}\left(1+\frac{|a-b|}{|-b+j-z_{0}|}\right)=C\prod_{\begin{subarray}{c}1\leqslant j\leqslant n\\ |j-k|\geqslant 2\end{subarray}}\left(1+\frac{|a-b|}{\left|j-k-\varepsilon\,\mathrm{e}^{\mathrm{i}\beta}\right|}\right)$
$\leqslant C\prod_{\begin{subarray}{c}1\leqslant j\leqslant n\\ |j-k|\geqslant 2\end{subarray}}\left(1+\frac{|a-b|}{|j-k|-4^{-1}}\right)\leqslant C\prod_{1\leqslant\ell\leqslant n}\left(1+\frac{|a-b|}{\ell}\right)^{2}$
$\leqslant C\cdot\exp\left(\sum_{1\leqslant\ell\leqslant n}\frac{2|a-b|}{\ell}\right)\leqslant C\left(n+1\right)^{2|a-b|},$
where the inequality $\sum_{j=1}^{n}\frac{1}{j}\leqslant 1+\log n$ is used.
Case 2. The proof is akin to that of Case 1, if one of the following conditions holds: (i) $\Re(z+b)\in[k+\varepsilon,k+1-\varepsilon]$ for some $k\in\mathbb{Z}\cap[1,n]$, (ii) $\Re(z+b)\leqslant-\varepsilon$, (iii) $\Re(z+b)\geqslant n+\varepsilon$. ∎
Now we establish explicit expansions of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ in different cases. The first gives expansions of ${}_{2}F_{2}[a,b-n;c,d-n;z]$ for large $z\in\mathbb{C}\setminus\mathbb{R}_{\geqslant 0}$.
Theorem 2.4.
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\geqslant 0}$ and $d,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\{\Re(a),\Re(b),\Re(d)\}$ and that the fractional parts of $w-\Re(a)$ and $w-\Re(b)$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is a small number. Then for fixed $n\in\mathbb{Z}_{\geqslant 0}$,
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};-z\right]=\frac{\Gamma(c)\Gamma(d-n)}{\Gamma(a)\Gamma(b-n)}\left\{S_{n}(z)+T_{n}(z)+R_{n,w}(z)\right\}$ (2.2)
as $z\to\infty$ such that $\left|\arg(z)\right|<\pi$ and $z$ is bounded way from the points $-b+k\,(k\in\mathbb{Z})$, where
$S_{n}(z)$ $=\sum_{k=0}^{\lfloor w-\Re(a)\rfloor}\frac{\Gamma(a+k)\Gamma(b-a-n-k)}{\Gamma(c-a-k)\Gamma(d-a-n-k)}\frac{\left(-1\right)^{k}}{k!}z^{-a-k},$
$T_{n}(z)$ $=\sum_{k=0}^{\lfloor w-\Re(b)\rfloor+n}\frac{\Gamma(b-n+k)\Gamma(a-b+n-k)}{\Gamma(d-b-k)\Gamma(c-b+n-k)}\frac{\left(-1\right)^{k}}{k!}z^{n-b-k},$
and
$R_{n,w}(z)=\mathcal{O}\left(\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left|z\right|^{-w}\right)+\mathcal{O}\left(\left(n+1\right)^{2|b-d|}\left|z\right|^{\Re(a+b-c-d)}\mathrm{e}^{-\Re(z)}\right).$
Proof.
Let $T$ be a positive number such that
$T>\max\{1,|\Im(a)|,|\Im(b)|,|\Im(c)|,|\Im(d)|\}.$
Denote $\theta=\arg(z)$. For $|\theta|<\pi$ and $n\in\mathbb{Z}_{\geqslant 0}$, define
$R_{n,w}(z)=\frac{1}{2\pi\mathrm{i}}\int_{\mathcal{C}}h_{n}(s)z^{s}\mathrm{d}s,\quad h_{n}(s)=\frac{\Gamma(a+s)\Gamma(b-n+s)}{\Gamma(c+s)\Gamma(d-n+s)}\Gamma(-s),$ (2.3)
where $\mathcal{C}$ is a negative-oriented loop that consists of the vertical line
$L^{v}:\quad s=-w+\mathrm{i}t,\quad|t|\leqslant T$
and the contours $L^{\pm}$ which pass to infinity in the directions ±$\pm\theta_{0}\,(0<\theta_{0}<\frac{\pi}{2})$. Further, $\mathcal{C}$ is taken to embrace all the poles of $\Gamma(-s)$ and the points
$s=-a-k\ (0\leqslant k\leqslant\lfloor w-\Re(a)\rfloor),\quad s=-b+n-k\ (0\leqslant k\leqslant\lfloor w-\Re(b)\rfloor+n),$ (2.4)
and $L^{\pm}$ are taken to be bounded away from the points shown in (2.4).
The convergence of the integral as $|s|\to\infty$ is controlled by the term $\exp\{-|s|\cos\theta_{0}\cdot\log|s|\}$ in the behaviour of the gamma function for large $|s|$. Use of Cauchy’s residue theorem implies that
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};-z\right]=\frac{\Gamma(c)\Gamma(d-n)}{\Gamma(a)\Gamma(b-n)}(S_{n}(z)+T_{n}(z)+R_{n,w}(z)).$
See [6, Section 2] for details. Divide $R_{n,w}(z)$ into two parts:
$R_{n,w}(z)=\frac{1}{2\pi\mathrm{i}}\int_{L^{v}}+\frac{1}{2\pi\mathrm{i}}\int_{L^{\pm}}=:R_{v}(z)+R_{\pm}(z).$
It remains to estimate the integrals $R_{v}(z)$ and $R_{\pm}(z)$, respectively.
Step 1. Estimate of $R_{v}(z)$. One may claim that for $|t|\leqslant T$,
$\left|\frac{\Gamma(a-w+\mathrm{i}t)}{\Gamma(c-w+\mathrm{i}t)}\right|=\mathcal{O}(1),\quad\left|\Gamma(-w-\mathrm{i}t)\right|=\mathcal{O}(1).$ (2.5)
Moreover,
$\frac{\sin(d-w+\mathrm{i}t)}{\sin(b-w+\mathrm{i}t)}=\frac{\mathrm{e}^{\mathrm{i}(d-w)-t}-\mathrm{e}^{-\mathrm{i}(d-w)+t}}{\mathrm{e}^{\mathrm{i}(b-w)-t}-\mathrm{e}^{-\mathrm{i}(b-w)+t}}=\mathcal{O}(1).$ (2.6)
If $|t|\leqslant\max\{1,|\Im(d)|\}$, apply [8, equations (5.6.6) and (5.6.7)] and (2.1) to yield
$g_{n}(t):=\left|\frac{\Gamma(1-d+w+n-\mathrm{i}t)}{\Gamma(1-b+w+n-\mathrm{i}t)}\right|\leqslant C\left(n+1\right)^{\Re(d-b)},\quad n\geqslant 0.$
If $|t|>\max\{1,|\Im(d)|\}$, take $r>0$ such that $w>r>\Re(d)$. Put $z_{1}=(1+w+n-r)-\mathrm{i}(t+\Im(d))$ and then
$g_{n}(t)=\left|\frac{\Gamma(z_{1}+r-\Re(d))}{\Gamma(z_{1}+r+\mathrm{i}\Im(d)-b)}\right|=:\left|\frac{\Gamma(z_{1}+a_{1})}{\Gamma(z_{1}+b_{1})}\right|.$
When $\Re(b_{1}-a_{1})\geqslant 1$, Lemma 2.2 shows that
$g_{n}(t)\leqslant\frac{\Gamma(\Re(d-b))}{\left|\Gamma(d-b)\right|}\mathrm{e}^{\frac{\pi}{2}|\Im(d-b)|}\left(|z_{1}|+(r-\Re(d))\cos\theta\right)^{\Re(b-d)}\leqslant C\left|z_{1}\right|^{\Re(b-d)}.$
When $\Re(b_{1}-a_{1})<1$, put $A=\lceil\Re(a_{1}-b_{1})\rceil$ and $a_{1}^{\prime}=a_{1}-A-1$. Then $\Re(b_{1}-a_{1}^{\prime})\geqslant 1$ and further,
$g_{n}(t)$ $=\left|\frac{\Gamma(z_{1}+a_{1}^{\prime})}{\Gamma(z_{1}+b_{1})}(z_{1}+a_{1}-A-1)(z_{1}+a_{1}-A)\cdots(z_{1}+a_{1}-1)\right|$
$\leqslant C\left|z_{1}\right|^{\Re(a_{1}^{\prime}-b_{1})}\left|(z_{1}+a_{1}-A-1)\cdots(z_{1}+a_{1}-1)\right|$
$\leqslant C\left|z_{1}\right|^{\Re(a_{1}-b_{1}-A-1)}\left|z_{1}\right|^{A+1}=C\left|z_{1}\right|^{\Re(d-b)}$
$=C\left(\left(1+w+n-r\right)^{2}+\left(t+\Im(d)\right)^{2}\right)^{\frac{1}{2}\Re(d-b)}$
$\leqslant C\cdot\max\left\{\left(n+1\right)^{\Re(d-b)},\left(|t|+1\right)^{\Re(d-b)}\right\}.$
In summary,
$g_{n}(t)\leqslant C\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left(|t|+1\right)^{\max\{0,\Re(d-b)\}}.$ (2.7)
Combining Euler’s reflection formula with the estimates (2.5)-(2.7) shows that for $s\in L^{v}$,
$|h_{n}(s)|\mathrm{e}^{-\theta t}$ $=\left|\frac{\Gamma(a-w+\mathrm{i}t)}{\Gamma(c-w+\mathrm{i}t)}\frac{\sin(d-w+\mathrm{i}t)}{\sin(b-w+\mathrm{i}t)}\frac{\Gamma(1-d+w+n-\mathrm{i}t)}{\Gamma(1-b+w+n-\mathrm{i}t)}\Gamma(-w-\mathrm{i}t)\right|\mathrm{e}^{-\theta t}$
$\leqslant C\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left(|t|+1\right)^{\max\{0,\Re(d-b)\}}\mathrm{e}^{-\theta t},$
Hence
$|R_{v}(z)|\leqslant\frac{1}{2\pi}\left|z\right|^{-w}\int_{-T}^{T}|h_{n}(s)|\mathrm{e}^{-\theta t}\mathrm{d}t\leqslant C\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left|z\right|^{-w}.$
Step 2. Estimate of $R_{\pm}(z)$. Due to [9, Lemma 2.1] and Euler’s reflection formula, we obtain that for $s\in L^{\pm}$,
$h_{0}(s)=\frac{\Gamma(a+s)\Gamma(b+s)}{\Gamma(c+d)\Gamma(d+s)\Gamma(1+s)}\frac{-\pi}{\sin(\pi s)}=\frac{-\pi\csc(\pi s)\rho_{0}(s)}{\Gamma(s+1-\alpha)}=\rho_{0}(s)\Gamma(-s+\alpha)\frac{\sin(\pi(s-\alpha))}{\sin(\pi s)},$
where $\alpha=a+b-c-d$ and $\rho_{0}(s)=\mathcal{O}(1)$ for $s\to\infty$ uniformly in $\left|\arg(s)\right|\leqslant\pi-\delta\,(\delta>0)$. Clearly,
$\left|\frac{\sin(\pi(s-\alpha))}{\sin(\pi s)}\right|=\mathcal{O}(1),\quad s\in L^{\pm}.$
Therefore, for $s\in L^{\pm}$, there are positive constants $K_{1}$ and $K_{2}$ independent of $n$ and $z$, such that when $|s|\geqslant K_{1}$,
$|\rho_{0}(s)|\leqslant K_{2},\quad\left|\frac{\sin(\pi(s-\alpha))}{\sin(\pi s)}\right|\leqslant K_{2}.$
In addition, it follows from Lemma 2.3 that for $s\in L^{\pm}$ and $n\in\mathbb{Z}_{\geqslant 0}$,
$\prod_{j=1}^{n}\left|\frac{-d+j-s}{-b+j-s}\right|=\mathcal{O}\left(\left(n+1\right)^{2|b-d|}\right).$
In view of the fact that
$h_{n}(s)=h_{0}(s)\prod_{j=1}^{n}\frac{d-j+s}{b-j+s}$
and the estimates above, we get
$\left|R_{\pm}(z)\right|$ $\leqslant C\left(n+1\right)^{2|b-d|}\left(\int_{s\in L^{\pm},|s|<K_{1}}|h_{0}(s)||ds|+\int_{s\in L^{\pm},|s|\geqslant K_{1}}|h_{0}(s)||ds|\right)$
$\leqslant C\left(n+1\right)^{2|b-d|}\left(K_{3}+K_{2}^{2}\int_{s\in L^{\pm},|s|\geqslant K_{1}}\left|\Gamma(-s+\alpha)z^{s}\right||ds|\right).$
According to [9, Lemma 2.8], the contours $L^{\pm}$ can be deformed to pass through the saddle point $s=-z$ and to be bounded away from the points shown in (2.4). Hence, the integral in the last line has the order of magnitude $\mathcal{O}(|z|^{\alpha}\mathrm{e}^{-\Re(z)})$.
The expansion (2.2) follows from the estimates of $R_{v}(z)$ and $R_{\pm}(z)$. ∎
The second provides expansions of ${}_{2}F_{2}[a,b+n;c,d+n;z]$ for large $z\in\mathbb{C}\setminus\mathbb{R}_{\geqslant 0}$.
Theorem 2.5.
Assume that $a,b\in\mathbb{C}$ and $c,d\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Then for $n\in\mathbb{Z}_{\geqslant\max\{0,-\Re(a),-\Re(b)\}}$,
${}_{2}F_{2}\left[\begin{matrix}a,b+n\\ c,d+n\end{matrix};-z\right]=\left(n+1\right)^{\max\{0,\Re(d-b)\}}\mathcal{O}(1)+\left(n+1\right)^{2|b-d|}\mathcal{O}\left(\left|z\right|^{\Re(a+b-c-d)}\mathrm{e}^{-\Re(z)}\right)$
as $z\to\infty$ such that $\left|\arg(z)\right|<\pi$ and $z$ is bounded way from the points $-b+k\,(k\in\mathbb{Z})$.
Proof.
It is sufficient to observe that
${}_{2}F_{2}\left[\begin{matrix}a,b+n\\ c,d+n\end{matrix};-z\right]=\frac{\Gamma(c)\Gamma(d+n)}{\Gamma(a)\Gamma(b+n)}R_{-n,0}(z),$
where $R_{-n,0}(z)$ is defined in (2.3). The result follows by repeating the proof of Theorem 2.4. ∎
The third establishes explicit upper bounds of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for $z\to+\infty$ and refines the rough bounds in [3, Lemma 3.3].
Theorem 2.6.
Let $p:=\Re(a-c)+\max\{0,\Re(b-d)\}$.
(i) Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $d\in\mathbb{C}\setminus\mathbb{Z}$. Then for $n\in\mathbb{Z}_{\geqslant 0}$,
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};z\right]=\mathcal{O}\left(\left(n+1\right)^{2|b-d|}z^{p}\mathrm{e}^{z}\right),\quad z\to+\infty.$
(ii) Assume that $a,b\in\mathbb{C}$ and $c,d\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Then for $n\in\mathbb{Z}_{\geqslant 0}$,
${}_{2}F_{2}\left[\begin{matrix}a,b+n\\ c,d+n\end{matrix};z\right]=\mathcal{O}\left(\left(n+1\right)^{2|b-d|}z^{p}\mathrm{e}^{z}\right),\quad z\to+\infty.$
Proof.
By follwing the proof of [3, Lemma 3.3], the results follow from Lemma 2.3. ∎
3. Proof of the main results
We are now in a position to prove our main results.
Proof of Theorem 1.2.
Recall the series representation (1.2). We have $V_{1}(x,y)=\mathcal{O}(1)$ since $\frac{y}{1-x}$ is bounded and then $V_{1}(x,y)$ is absolutely convergent. Hence the main contribution of $\Psi_{1}$ comes from $V_{2}(x,y)$. Now divide $V_{2}(x,y)$ into two parts:
$V_{2}(x,y)=\sum_{n=0}^{N-1}+\sum_{n=N}^{\infty}=:V_{2}^{(L)}(x,y)+V_{2}^{(R)}(x,y),$ (3.1)
where $N$ is a positive integer to be determined.
For $10\leqslant n\leqslant N-$ 1, recall the exponential expansion of ${}_{2}F_{2}$ [6, equation (5.8)]
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]=\frac{\Gamma(c)\Gamma(d-n)}{\Gamma(a)\Gamma(b-n)}\mathrm{e}^{y}\sum_{k=0}^{N-1}c_{k,n}y^{a+b-c-d-k}+\mathcal{O}\left(y^{\Re(a+b-c-d)-N}\mathrm{e}^{y}\right),$ (3.2)
as $y\to+\infty$, where the coefficients $c_{k,n}$ are given by [10, equation (12)]
$c_{k,n}:=\frac{\left(c+d-a-b\right)_{k}\left(n+1-b\right)_{k}}{k!}{}_{3}F_{2}\left[\begin{matrix}-k,c-a,d-a-n\\ c+d-a-b,b-n-k\end{matrix};1\right].$ (3.3)
Inserting (3.2) into (3.1), we can infer that
$V_{2}^{(L)}(x,y)=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{n=0}^{N-1}\frac{\left(b\right)_{n}}{n!}\left(1-x\right)^{-n}\sum_{k=0}^{N-1}c_{k,n}^{*}y^{-k}+\mathcal{O}\left(y^{\Re(a+b-c-d)-N}\mathrm{e}^{y}\right),$
where
$c_{k,n}^{*}=\frac{\left(b-a+c^{\prime}\right)_{k}\left(n+b-a+1\right)_{k}}{k!}{}_{3}F_{2}\left[\begin{matrix}-k,-n,c+c^{\prime}-a-1\\ b-a+c^{\prime},a-b-n-k\end{matrix};1\right].$
Hence
$V_{2}^{(L)}(x,y)$ $=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}y^{-k}\sum_{j=0}^{k}c_{k-j,j}^{*}\frac{\left(b\right)_{j}}{j!}\Big{(}\frac{y}{1-x}\Big{)}^{j}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)\left(\frac{1-x}{y}\right)^{b}y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-N}\mathrm{e}^{y}\right),$ (3.4)
where $a_{k}(x,y)$ is given by (1.3).
For $n\geqslant N$, a combination of (2.1) and Theorem 2.6 yields
$\left|V_{2}^{(R)}(x,y)\right|$ $\leqslant Cy^{p}\mathrm{e}^{y}\sum_{n=N}^{\infty}\left(n+1\right)^{q}\left|1-x\right|^{-n}\leqslant Cy^{p}\mathrm{e}^{y}\sum_{n=N}^{\infty}\left(n+1\right)^{q}\gamma_{2}^{n}y^{-n}$
$=C\gamma_{2}^{N}y^{p-N}\mathrm{e}^{y}\sum_{k=0}^{\infty}\left(k+N+1\right)^{q}\gamma_{2}^{k}y^{-k}=\mathcal{O}\left(y^{p-N}\mathrm{e}^{y}\right),$ (3.5)
where $p=\Re(a-c-c^{\prime}+1)+\max\{0,\Re(c-b-1)\}$ and $q=\Re(c)-2+2|b-c+1|$.
For any given positive integer $M$, take $N=M+\lceil p-\Re(a-b-c^{\prime})\rceil$. Since $p\geqslant\Re(a-b-c^{\prime})$, one has $N\geqslant M$ and $\Re(a-b-c^{\prime})-M\geqslant p-N$. Then it follows from (3.4) and (3.5) that
$\left(1-x\right)^{-b}V_{2}(x,y)$ $=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{p-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-M}\mathrm{e}^{y}\right)+\mathcal{O}\left(y^{p-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-M}\mathrm{e}^{y}\right),$
which completes the proof. ∎
Proof of Theorem 1.3.
Estimates of error terms here are similar to those of Theorem 1.2, so we omit the details and just extract the expansions.
Let us start with a convergent series
$F(x,y):=\frac{\Gamma(a)\Gamma(b)}{\Gamma(c)\Gamma(d)}\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}}{}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]\left(1-x\right)^{-n}.$
Recall the asymptotic expansion of ${}_{2}F_{2}(y)$ [6, equation (5.8)]: for $0\leqslant n\leqslant N$,
$\begin{split}{}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]={}&\frac{\Gamma(c)\Gamma(d)}{\Gamma(a)\Gamma(b)}\frac{\left(a_{3}\right)_{n}}{\left(b_{3}\right)_{n}}\bigg{\{}S_{n}(-y)+T_{n}(-y)+\mathrm{e}^{y}\sum_{k=0}^{N-1}c_{k,n}y^{a+b-c-d-k}\\ &+\mathcal{O}\left(\left|y\right|^{-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a+b-c-d)-N}\mathrm{e}^{\Re(y)}\right)\bigg{\}},\end{split}$ (3.6)
where $a_{3}=1-b,\,b_{3}=1-d$, $S_{n}(-y)$ and $T_{n}(-y)$ are defined in Theorem 2.4, and the coefficients $c_{k,n}$ are given by (3.3). For $n\geqslant N+1$, recall the expansion (2.2). Use of (2.1) and the definition of $F(x,y)$ then gives
$F(x,y)=S_{w}(x,y)+T_{w}(x,y)+E_{N}(x,y)+\mathcal{O}\left(\left|y\right|^{-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a+b-c-d)-N}\mathrm{e}^{\Re(y)}\right),$ (3.7)
where
$S_{w}(x,y)$ $:=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}}S_{n}(-y)\left(1-x\right)^{-n},$
$T_{w}(x,y)$ $:=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}}T_{n}(-y)\left(1-x\right)^{-n},$
and
$E_{N}(x,y):=y^{a+b-c-d}\mathrm{e}^{y}\sum_{k=0}^{N-1}y^{-k}\sum_{j=0}^{k}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}c_{k-j,j}\left(\frac{y}{1-x}\right)^{j}.$
Dividing $T_{w}(x,y)$ into two parts:
$T_{w}(x,y)=\sum_{n=0}^{\infty}\sum_{k=0}^{n}+\sum_{n=0}^{\infty}\sum_{k=n+1}^{\lfloor w-\Re(b)\rfloor+n}=:T_{w}^{-}(x,y)+T_{w}^{+}(x,y),$
applying the transformation [7, p. 100, equation (1)]
$\sum_{n=0}^{\infty}\sum_{k=0}^{n}A(k,n-k)=\sum_{n=0}^{\infty}\sum_{k=0}^{\infty}A(k,n)$
to $T_{w}^{-}(x,y)$ and using the series manipulation technique, we have
$S_{w}(x,y)={}$ $\frac{\Gamma(a)\Gamma(b-a)}{\Gamma(c-a)\Gamma(d-a)}\sum_{k=0}^{\lfloor w-\Re(a)\rfloor}\frac{\left(a\right)_{k}\left(a-c+1\right)_{k}\left(a-d+1\right)_{k}}{\left(1\right)_{k}\left(a-b+1\right)_{k}}$
$\qquad\cdot\left(-y\right)^{-a-k}{}_{5}F_{4}\left[\begin{matrix}a_{1},a_{2},a_{3},1,a-d+k+1\\ b_{1},b_{2},b_{3},a-b+k+1\end{matrix};\frac{1}{1-x}\right]$
[…]
$T_{w}^{+}(x,y)={}$ $\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\sum_{k=1}^{\lfloor w-\Re(b)\rfloor}\frac{\left(b\right)_{k}\left(b-c+1\right)_{k}\left(b-d+1\right)_{k}}{\left(1\right)_{k}\left(b-a+1\right)_{k}}$
$\qquad\cdot\left(-y\right)^{-b-k}{}_{5}F_{4}\left[\begin{matrix}a_{1},a_{2},a_{3},1,b-d+k+1\\ b_{1},b_{2},b_{3},k+1\end{matrix};\frac{1}{1-x}\right]$
[…]
$T_{w}^{-}(x,y)={}$ $\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\sum_{k=0}^{\infty}\frac{\left(a_{1}\right)_{k}\left(a_{2}\right)_{k}\left(a_{3}\right)_{k}\left(b-d+1\right)_{k}}{\left(b_{1}\right)_{k}\left(b_{2}\right)_{k}\left(b_{3}\right)_{k}\left(1\right)_{k}}t_{k}^{-}(x,y)\left(-y\right)^{-b-k},$
where
$s_{k}(x,y)$ $=\sum_{j=0}^{k}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}\frac{\left(a\right)_{k-j}\left(a-c+1\right)_{k-j}}{\left(1\right)_{k-j}}\left(\frac{y}{x-1}\right)^{j},$
$t_{k}^{+}(x,y)$ $=\sum_{j=0}^{k-1}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}\frac{\left(b\right)_{k-j}\left(b-c+1\right)_{k-j}}{\left(b-a+1\right)_{k-j}}\left(\frac{y}{x-1}\right)^{j},$
$t_{k}^{-}(x,y)$ $={}_{4}F_{5}\left[\begin{matrix}a_{1}+k,a_{2}+k,a_{3}+k,a-b\\ b_{1}+k,b_{2}+k,b_{3}+k,1-b,c-b\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{k}.$
The asymptotic expansion of $\Psi_{1}$ follows from (1.2) and the expansion (3.7). ∎
4. Concluding remarks
We have established the complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments. Our derivation is based on the estimates of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for large $z$ (see Theorems 2.42.6), but the exponential expansions are not included.
We conjecture that the more accurate asymptotic expansion (3.6) is valid for $n\in\mathbb{Z}$ with the remainder terms multiplied by $n^{\lambda}$ for some $\lambda>0$. If this is the case, we may deal with some revelant problems:
  • (i)
    Asymptotics of $\Psi_{1}[a,b;c,c^{\prime};x,y]$ for large $y$: This problem was partially solved in [4, Section 3.1]. We guess that the series representations (1.2) and [3, equation (3.4)] offer the complete asymptotic expansions.
  • (ii)
    Asymptotics of $\Phi_{3}[b;c;w,z]$ for large $z$ : One may deduce the asymptotic behavior via the series representation [2, equation (1.13)]
    $\Phi_{3}[b;c;w,z]=\sum_{m=0}^{\infty}\frac{\left(b\right)_{m}}{\left(c\right)_{m}}{}_{1}F_{1}\left[\begin{matrix}-m\\ 1-b-m\end{matrix};\frac{z}{w}\right]\frac{w^{m}}{m!}.$
  • (iii)
    Asymptotics of $\Phi_{1},\Phi_{2},\Xi_{1}$ for large $y$: See the series representations shown in [5, p. 76].
Braaksma’s work [1] seems helpful to examine this conjecture.
Finally, we are not aware right now of any application of our Theorems 1.2 and 1.3.
Acknowledgements
The author thanks Prof. Nico M. Temme for bringing Braaksma’s work to his attention.