### v4

# Complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments

Abstract
In our recent work [SIGMA 20, 074 (2024)], the leading behaviour of the Humbert function $\Psi_{1}[a,b;c,c^{\prime};x,y]$ when $x\to\infty$ and $y\to+\infty$ has been derived in a direct and simple manner. In this paper, we obtain the complete asymptotics of $\Psi_{1}$ in the general case $x,y\to\infty$ along a new path. Indeed, our proof is based on a sharp estimate on ${}_{2}F_{2}[a,b-n;c,d-n;z]$, which is valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$ and large $z$.
Mathematics Subject Classification: 33C70, 41A60, 30E15.
Keywords: Humbert functions, asymptotic expansions.
1.  Introduction
The confluent hypergeometric function of two variables $\Psi_{1}$, introduced by Humbert [7, p. 75], is defined by
$\Psi_{1}[a,b;c,c^{\prime};x,y]:=\sum_{m,n=0}^{\infty}\frac{(a)_{m+n}(b)_{m}}{(c)_{m}(c^{\prime})_{n}}\frac{x^{m}}{m!}\frac{y^{n}}{n!},\qquad|x|<1,|y|<\infty,$
where $a,b\in\mathbb{C}$ and $c,c^{\prime}\notin\mathbb{Z}_{\leqslant 0}$. It is known that $\Psi_{1}$ has an extension to the region (see [4, 5])
$\mathbb{D}_{\Psi_{1}}:=\bigl{\{}(x,y)\in\mathbb{C}^{2}\colon x\neq 1,\,\left|\arg(1-x)\right|<\pi,\,|y|<\infty\bigr{\}}.$ (1.1)
By using inverse Laplace transformations of the Humbert functions $\Phi_{2}$, $\Phi_{3}$ and $\Xi_{2}$, Wald and Henkel [13] derived the leading-order behaviour of these functions when two variables become simultaneously large. But they failed to interpret the integral for $\Psi_{1}$ [13, Equation (2.4b)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{1}{\Gamma(a)}\int_{0}^{\infty}\mathrm{e}^{-u}u^{a-1}\,_{1}F_{1}\left[\begin{matrix}b\\ c\end{matrix};xu\right]{}_{0}F_{1}\left[\begin{matrix}-\\ c^{\prime}\end{matrix};yu\right]\mathrm{d}u$ (1.2)
as a convolution and so could not establish the asymptotics of $\Psi_{1}$.
Recently, we [4] partially answered Wald and Henkel’s problem by deriving the leading behaviour of $\Psi_{1}$ under the condition
$x\to\infty,y\to+\infty,\quad\left|\arg(1-x)\right|<\pi,\quad 0<\gamma_{1}\leqslant\frac{y}{\left|1-x\right|}\leqslant\gamma_{2}<\infty.$ (1.3)
Our starting point is the following expression for $\Psi_{1}$.
Theorem 1.1 ([4, Theorem 3.1]).
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c\in\mathbb{C}\setminus\mathbb{Z}$. Then
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}\left(1-x\right)^{-a}V_{1}(x,y)+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)}\left(1-x\right)^{-b}V_{2}(x,y)$ (1.4)
holds for $|{\arg}(1-x)|<\pi$, $|x-1|>1$ and $|y|<\infty$, where
$V_{1}(x,y):=\sum_{n=0}^{\infty}\frac{\left(a\right)_{n}\left(c-b\right)_{n}}{\left(a-b+1\right)_{n}}\,_{2}F_{2}\left[\begin{matrix}a-c+1,a+n\\ c^{\prime},a-b+1+n\end{matrix};\frac{y}{1-x}\right]\frac{\left(1-x\right)^{-n}}{n!},$
$V_{2}(x,y):=\sum_{n=0}^{\infty}\frac{\left(b\right)_{n}\left(c-a\right)_{n}}{\left(b-a+1\right)_{n}}\,_{2}F_{2}\left[\begin{matrix}a-c+1,a-b-n\\ c^{\prime},a-c+1-n\end{matrix};y\right]\frac{\left(1-x\right)^{-n}}{n!}.$
In this paper, we shall give detailed asymptotic analyses of $\Psi_{1}$ when $x\to\infty$ and $y\to\infty$. Our first result provides the full asymptotic expansion of $\Psi_{1}$ under the condition (1.3), which covers our previous result [4, Theorem 3.6].
Theorem 1.2.
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c\in\mathbb{C}\setminus\mathbb{Z}$. Then, under the condition (1.3),
$\Psi_{1}[a,b;c,c^{\prime};x,y]\sim\frac{\Gamma(c)\Gamma(c^{\prime})}{\Gamma(a)\Gamma(c-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{\infty}a_{k}(x,y)y^{-k},$
where $a_{0}(x,y)=\bigl{(}\frac{y}{1-x}\bigr{)}^{b}$ and in general, for any $k\in\mathbb{Z}_{\geqslant 0}$,
$a_{k}(x,y)=\sum_{j=0}^{k}\frac{\left(b\right)_{j}\left(b-a+c^{\prime}\right)_{k-j}\left(j+b-a+1\right)_{k-j}}{j!\left(k-j\right)!}\,_{3}F_{2}\left[\begin{matrix}-j,j-k,c+c^{\prime}-a-1\\ b-a+c^{\prime},a-b-k\end{matrix};1\right]\left(\frac{y}{1-x}\right)^{b+j}.$ (1.5)
Our second result gives the full asymptotics of $\Psi_{1}$ for two large arguments in the remaining cases. These two results give a complete answer to Wald and Henkel’s problem.
Theorem 1.3.
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c,b-c\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\left\{\Re(a-b)+1,\Re(a-c)+2\right\}$ and that the fractional parts of $w-\Re(a-b)$ and $w-\Re(a-c)-1$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is small.
Then, under the condition
$x\to\infty,y\to\infty,\quad\left|\arg(1-x)\right|<\pi,\quad\left|\arg(-y)\right|<\pi,\quad 0<\gamma_{1}\leqslant\left|\frac{y}{1-x}\right|\leqslant\gamma_{2}<\infty$ (1.6)
and under the restriction that $y$ is bounded away from the points $b-a+k\,(k\in\mathbb{Z})$, the function $\Psi_{1}\equiv\Psi_{1}[a,b;c,c^{\prime};x,y]$ admits the asymptotic expansion
$\Psi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}(x,y)+\frac{\Gamma(c)\Gamma(c^{\prime})\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)\Gamma(b-a+c^{\prime})}A_{2}(x,y)+\frac{\Gamma(c)\Gamma(c^{\prime})}{\Gamma(a)\Gamma(c-b)}A_{3}(x,y)$
$+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-2b-c^{\prime})-N}\mathrm{e}^{\Re(y)}\right),$
where
$A_{1}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(a\right)_{k}\left(c-b\right)_{k}}{\left(a-b+1\right)_{k}k!}\,_{2}F_{2}\left[\begin{matrix}a-c+1,a+k\\ c^{\prime},a-b+1+k\end{matrix};\frac{y}{1-x}\right]\left(1-x\right)^{-a-k},$
$A_{2}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(a-b\right)_{k}\left(a-b-c^{\prime}+1\right)_{k}}{k!}\,_{2}F_{2}\left[\begin{matrix}b,b-c+1-k\\ b-a+1-k,b-a+c^{\prime}-k\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{b}\left(-y\right)^{-a-k},$
$A_{3}(x,y)={}$ $y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k},$
with $M=\lfloor w+\Re(b-a)\rfloor\geqslant 1$, $N$ being any positive integer and $a_{k}(x,y)$ given by (1.5).
The paper is organized as follows. In Section 2, we demonstrate some estimates of the generalized hypergeometric functions ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ when $z$ goes all the way to infinity, which are valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$. In Section 3, we use the tools in Section 2 to prove our main results. The paper concludes with further remarks in Section 4.
Notation. The number $C$ generically denotes a positive constant independent of the summation index $n$ and the variable $z$. By $f(n,z)=\mathcal{O}(a_{n}g(z))\,(z\in\Omega)$, we mean that there exists a constant $K>0$ independent of $n$ and $z$ such that
$|f(n,z)|\leqslant K|a_{n}||g(z)|,\quad n\in\mathbb{Z}_{\geqslant 0},\,z\in\Omega.$
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by [10, Equation (16.2.1)]
${}_{p}F_{q}\left[\begin{matrix}a_{1},\cdots,a_{p}\\ b_{1},\cdots,b_{q}\end{matrix};z\right]\equiv{}_{p}F_{q}[a_{1},\cdots,a_{p};b_{1},\cdots,b_{q};z]:=\sum_{n=0}^{\infty}\frac{(a_{1})_{n}\cdots(a_{p})_{n}}{(b_{1})_{n}\cdots(b_{q})_{n}}\frac{z^{n}}{n!},$ (1.7)
where $a_{1},\cdots,a_{p}\in\mathbb{C}$ and $b_{1},\cdots,b_{q}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Empty products and sums are taken as $1$ and 00, respectively.
2.  Auxiliary results
In this section, we prove some auxiliary results, including three lemmas about explicit bounds for the ratio of two gamma functions: (i) the first is beneficial to bound the tail series; (ii) the second generalizes the simple bound of gamma ratio [10, Equation (5.6.8)]; (iii) the third guarantees the uniformity of explicit bounds for ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ obtained in the ensuing theorems.
Lemma 2.1 ([4, Lemma 2.1]).
If $a\in\mathbb{C}$ and $b\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$, then
$\left|\frac{\left(a\right)_{n}}{\left(b\right)_{n}}\right|\leqslant C\left(n+1\right)^{\Re(a-b)},\qquad n\in\mathbb{Z}_{\geqslant 0}.$ (2.1)
Lemma 2.2.
If $\Re(b)>\Re(a)\geqslant 0$ and $\Re(z)>|\Im(a)|$, then
$\left|\frac{\Gamma(z+a)}{\Gamma(z+b)}\right|\leqslant\frac{\Gamma(\Re(b-a))}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\left(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta\right)^{\Re(a-b)},$
where $\theta=\arg(z)\in(-\frac{\pi}{2},\frac{\pi}{2})$.
Proof..
Recall the integral representation for the quotient of two gamma functions [11, p. 33]: for $b>a\geqslant 0$ and $\Re(z)>0$,
$\frac{\Gamma(z+a)}{\Gamma(z+b)}=\frac{\mathrm{e}^{-\mathrm{i}\theta}}{\Gamma(b-a)}\int_{0}^{\infty}\mathrm{e}^{-|z|u-au\mathrm{e}^{-\mathrm{i}\theta}}\left(1-\mathrm{e}^{-u\mathrm{e}^{-\mathrm{i}\theta}}\right)^{b-a-1}\mathrm{d}u.$
By analytic continuation, it is valid for $\Re(b)>\Re(a)\geqslant 0$ and $\Re(z)>|\Im(a)|$. Note that
$1-\mathrm{e}^{-u\mathrm{e}^{-\mathrm{i}\theta}}=1-\mathrm{e}^{-u\cos\theta}\cos(u\sin\theta)-\mathrm{i}\cdot\mathrm{e}^{-u\cos\theta}\sin(u\sin\theta)$
lies in the right half-plane and $|1-\mathrm{e}^{-\zeta}|\leqslant|\zeta|$ when $\left|\arg(\zeta)\right|\leqslant\frac{\pi}{2}$. Hence
$\left|\frac{\Gamma(z+a)}{\Gamma(z+b)}\right|$ $\leqslant\frac{1}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\int_{0}^{\infty}\mathrm{e}^{-(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta)u}u^{\Re(b-a)-1}\mathrm{d}u$
$=\frac{\Gamma(\Re(b-a))}{|\Gamma(b-a)|}\mathrm{e}^{\frac{\pi}{2}|\Im(a-b)|}\left(|z|+\Re(a)\cos\theta+\Im(a)\sin\theta\right)^{\Re(a-b)},$
where the identity
$\int_{0}^{\infty}u^{\alpha-1}\mathrm{e}^{-\lambda u}\mathrm{d}u=\frac{\Gamma(\alpha)}{\lambda^{\alpha}}~{}~{}~{}(\min\{\Re(\alpha),\Re(\lambda)\}>0)$
is used. ∎
Lemma 2.3.
If $a,b\in\mathbb{C}$, then
$\prod_{j=1}^{n}\left|\frac{-a+j-z}{-b+j-z}\right|\leqslant C\left(n+1\right)^{2|a-b|},\quad n\in\mathbb{Z}_{\geqslant 0}$
holds for $z$ bounded away from the points $-b+1,\cdots,-b+n$.
Proof..
Suppose that $n\geqslant 1$ and $z\in\bigcup_{j=1}^{n}\big{\{}|z+b-j|\geqslant\varepsilon\big{\}}$ for some $\varepsilon\in\big{(}0,\frac{1}{4}\big{)}$. Denote
$P_{n}(z):=\prod_{j=1}^{n}\left|\frac{-a+j-z}{-b+j-z}\right|.$
Case 1. If $\Re(z+b)\in(k-\varepsilon,k+\varepsilon)$ for some $k\in\mathbb{Z}\cap[1,n]$, there is a point $z_{0}=-b+k+\varepsilon\,\mathrm{e}^{\mathrm{i}\beta}$ with $\Re(z)=\Re(z_{0})$, such that for $j\in\mathbb{Z}\cap[1,n]$,
$\left|-b+j-z\right|\geqslant\left|-b+j-z_{0}\right|.$
Therefore,
$P_{n}(z)$ $\leqslant\prod_{j=1}^{n}\left(1+\frac{|a-b|}{\left|-b+j-z_{0}\right|}\right)=C\prod_{\begin{subarray}{c}1\leqslant j\leqslant n\\ |j-k|\geqslant 2\end{subarray}}\left(1+\frac{|a-b|}{\left|j-k-\varepsilon\,\mathrm{e}^{\mathrm{i}\beta}\right|}\right)$
$\leqslant C\prod_{\begin{subarray}{c}1\leqslant j\leqslant n\\ |j-k|\geqslant 2\end{subarray}}\left(1+\frac{|a-b|}{|j-k|-4^{-1}}\right)\leqslant C\prod_{1\leqslant\ell\leqslant n}\left(1+\frac{|a-b|}{\ell}\right)^{2}$
$\leqslant C\cdot\exp\left(\sum_{1\leqslant\ell\leqslant n}\frac{2|a-b|}{\ell}\right)\leqslant C\left(n+1\right)^{2|a-b|},$
where the inequality $\sum_{j=1}^{n}\frac{1}{j}\leqslant 1+\log n$ is used.
Case 2. The proof is akin to that of Case 1, if one of the following conditions holds: (i) $\Re(z+b)\in[k+\varepsilon,k+1-\varepsilon]$ for some $k\in\mathbb{Z}\cap[1,n]$, (ii) $\Re(z+b)\leqslant-\varepsilon$, (iii) $\Re(z+b)\geqslant n+\varepsilon$. ∎
Now we establish explicit expansions of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ in different cases. The first gives expansions of ${}_{2}F_{2}[a,b-n;c,d-n;z]$ for large $z\in\mathbb{C}\setminus\mathbb{R}_{\geqslant 0}$.
Theorem 2.4.
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\geqslant 0}$ and $d,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be such that $w>\max\{\Re(a),\Re(b),\Re(d)\}$ and that the fractional parts of $w-\Re(a)$ and $w-\Re(b)$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is a small number. Then for any $n\in\mathbb{Z}_{\geqslant 0}$,
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};-z\right]=\frac{\Gamma(c)\Gamma(d-n)}{\Gamma(a)\Gamma(b-n)}\left\{S_{n}(z)+T_{n}(z)+R_{n,w}(z)\right\}$ (2.2)
as $z\to\infty$ such that $\left|\arg(z)\right|<\pi$ and $z$ is bounded way from the points $-b+k\,(k\in\mathbb{Z})$, where
$S_{n}(z)$ $=\sum_{k=0}^{\lfloor w-\Re(a)\rfloor}\frac{\Gamma(a+k)\Gamma(b-a-n-k)}{\Gamma(c-a-k)\Gamma(d-a-n-k)}\frac{\left(-1\right)^{k}}{k!}z^{-a-k},$
$T_{n}(z)$ $=\sum_{k=0}^{\lfloor w-\Re(b)\rfloor+n}\frac{\Gamma(b-n+k)\Gamma(a-b+n-k)}{\Gamma(d-b-k)\Gamma(c-b+n-k)}\frac{\left(-1\right)^{k}}{k!}z^{n-b-k},$
and
$R_{n,w}(z)=\mathcal{O}\left(\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left|z\right|^{-w}\right)+\mathcal{O}\left(\left(n+1\right)^{2|b-d|}\left|z\right|^{\Re(a+b-c-d)}\mathrm{e}^{-\Re(z)}\right).$
Proof..
Let $T$ be a positive number such that
$T>\max\big{\{}1,|\Im(a)|,|\Im(b)|,|\Im(c)|,|\Im(d)|\big{\}}.$
Denote $\theta=\arg(z)$. For $|\theta|<\pi$ and $n\in\mathbb{Z}_{\geqslant 0}$, define
$R_{n,w}(z)=\frac{1}{2\pi\mathrm{i}}\int_{\mathcal{C}}h_{n}(s)z^{s}\mathrm{d}s,\quad h_{n}(s)=\frac{\Gamma(a+s)\Gamma(b-n+s)}{\Gamma(c+s)\Gamma(d-n+s)}\Gamma(-s),$ (2.3)
where $\mathcal{C}$ is a negative-oriented loop that consists of the vertical line
$L^{v}:\quad s=-w+\mathrm{i}t,\quad|t|\leqslant T$
and the contours $L^{\pm}$ which pass to infinity in the directions ±$\pm\theta_{0}$ $\big{(}0<\theta_{0}<\frac{\pi}{2}\big{)}$. Further, $\mathcal{C}$ is taken to embrace all the poles of $\Gamma(-s)$ and the points
$s=-a-k\ (0\leqslant k\leqslant\lfloor w-\Re(a)\rfloor),\quad s=-b+n-k\ (0\leqslant k\leqslant\lfloor w-\Re(b)\rfloor+n),$ (2.4)
and $L^{\pm}$ are taken to be bounded away from the points shown in (2.4).
The convergence of the integral as $|s|\to\infty$ is guaranteed by the term $\exp\{-|s|\cos\theta_{0}\cdot\log|s|\}$ in the behaviour of the gamma function for large $|s|$. The use of Cauchy’s residue theorem implies that
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};-z\right]=\frac{\Gamma(c)\Gamma(d-n)}{\Gamma(a)\Gamma(b-n)}(S_{n}(z)+T_{n}(z)+R_{n,w}(z)).$
See [8, Section 2] for details. Divide $R_{n,w}(z)$ into two parts:
$R_{n,w}(z)=\frac{1}{2\pi\mathrm{i}}\int_{L^{v}}+\frac{1}{2\pi\mathrm{i}}\int_{L^{\pm}}=:R_{v}(z)+R_{\pm}(z).$
It remains to estimate the integrals $R_{v}(z)$ and $R_{\pm}(z)$, respectively.
Step 1. Estimate of $R_{v}(z)$. One may claim that for $|t|\leqslant T$,
$\left|\frac{\Gamma(a-w+\mathrm{i}t)}{\Gamma(c-w+\mathrm{i}t)}\Gamma(-w-\mathrm{i}t)\right|=\mathcal{O}(1).$ (2.5)
Moreover,
$\frac{\sin(d-w+\mathrm{i}t)}{\sin(b-w+\mathrm{i}t)}=\frac{\mathrm{e}^{\mathrm{i}(d-w)-t}-\mathrm{e}^{-\mathrm{i}(d-w)+t}}{\mathrm{e}^{\mathrm{i}(b-w)-t}-\mathrm{e}^{-\mathrm{i}(b-w)+t}}=\mathcal{O}(1).$ (2.6)
If $|t|\leqslant\max\{1,|\Im(d)|\}$, apply [10, Equations (5.6.6) and (5.6.7)] and (2.1) to yield
$g_{n}(t):=\left|\frac{\Gamma(1-d+w+n-\mathrm{i}t)}{\Gamma(1-b+w+n-\mathrm{i}t)}\right|\leqslant C\left(n+1\right)^{\Re(d-b)},\quad n\geqslant 0.$
If $|t|>\max\{1,|\Im(d)|\}$, take $r>0$ such that $w>r>\Re(d)$. Set $z_{1}=(1+w+n-r)-\mathrm{i}(t+\Im(d))$ and then
$g_{n}(t)=\left|\frac{\Gamma(z_{1}+r-\Re(d))}{\Gamma(z_{1}+r+\mathrm{i}\Im(d)-b)}\right|=:\left|\frac{\Gamma(z_{1}+a_{1})}{\Gamma(z_{1}+b_{1})}\right|.$
When $\Re(b_{1}-a_{1})=\Re(d-b)\geqslant 1$, Lemma 2.2 shows that
$g_{n}(t)\leqslant\frac{\Gamma(\Re(d-b))}{\left|\Gamma(d-b)\right|}\mathrm{e}^{\frac{\pi}{2}|\Im(d-b)|}\left(|z_{1}|+(r-\Re(d))\cos\theta\right)^{\Re(b-d)}\leqslant C\left|z_{1}\right|^{\Re(b-d)}.$
When $\Re(b_{1}-a_{1})<1$, let $A=\lceil\Re(a_{1}-b_{1})\rceil$ and $a_{1}^{\prime}=a_{1}-A-1$. Then $\Re(b_{1}-a_{1}^{\prime})\geqslant 1$ and further,
$g_{n}(t)$ $=\left|\frac{\Gamma(z_{1}+a_{1}^{\prime})}{\Gamma(z_{1}+b_{1})}(z_{1}+a_{1}-A-1)(z_{1}+a_{1}-A)\cdots(z_{1}+a_{1}-1)\right|$
$\leqslant C\left|z_{1}\right|^{\Re(a_{1}^{\prime}-b_{1})}\left|(z_{1}+a_{1}-A-1)\cdots(z_{1}+a_{1}-1)\right|$
$\leqslant C\left|z_{1}\right|^{\Re(a_{1}-b_{1}-A-1)}\left|z_{1}\right|^{A+1}=C\left|z_{1}\right|^{\Re(d-b)}$
$=C\left(\left(1+w+n-r\right)^{2}+\left(t+\Im(d)\right)^{2}\right)^{\frac{1}{2}\Re(d-b)}$
$\leqslant C\cdot\max\left\{\left(n+1\right)^{\Re(d-b)},\left(|t|+1\right)^{\Re(d-b)}\right\}.$
In summary,
$g_{n}(t)\leqslant C\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left(|t|+1\right)^{\max\{0,\Re(d-b)\}}.$ (2.7)
Combining Euler’s reflection formula with the estimates (2.5)-(2.7) shows that for $s\in L^{v}$,
$|h_{n}(s)|\mathrm{e}^{-\theta t}$ $=\left|\frac{\Gamma(a-w+\mathrm{i}t)}{\Gamma(c-w+\mathrm{i}t)}\frac{\sin(d-w+\mathrm{i}t)}{\sin(b-w+\mathrm{i}t)}\frac{\Gamma(1-d+w+n-\mathrm{i}t)}{\Gamma(1-b+w+n-\mathrm{i}t)}\Gamma(-w-\mathrm{i}t)\right|\mathrm{e}^{-\theta t}$
$\leqslant C\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left(|t|+1\right)^{\max\{0,\Re(d-b)\}}\mathrm{e}^{-\theta t},$
Hence
$|R_{v}(z)|\leqslant\frac{1}{2\pi}\left|z\right|^{-w}\int_{-T}^{T}|h_{n}(s)|\mathrm{e}^{-\theta t}\mathrm{d}t\leqslant C\left(n+1\right)^{\max\{0,\Re(d-b)\}}\left|z\right|^{-w}.$
Step 2. Estimate of $R_{\pm}(z)$. Due to [11, Lemma 2.1] and Euler’s reflection formula, we obtain that for $s\in L^{\pm}$,
$h_{0}(s)$ $=\frac{\Gamma(1-c-s)\Gamma(1-d-s)}{\Gamma(1-a-s)\Gamma(1-b-s)}\Gamma(-s)\cdot\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))}$
$=\rho_{0}(-s)\Gamma(-s+\alpha)\cdot\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))},$
where $\alpha=a+b-c-d$ and $\rho_{0}(s)=\mathcal{O}(1)$ for $s\to\infty$ uniformly in $\left|\arg(s)\right|\leqslant\pi-\delta\,(\delta>0)$. Clearly,
$\left|\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))}\right|=\mathcal{O}(1),\quad s\in L^{\pm}.$
Therefore, for $s\in L^{\pm}$, there are positive constants $K_{1}$ and $K_{2}$ independent of $n$ and $z$, such that when $|s|\geqslant K_{1}$ then
$|\rho_{0}(-s)|\leqslant K_{2},\quad\left|\frac{\sin(\pi(c+s))\sin(\pi(d+s))}{\sin(\pi(a+s))\sin(\pi(b+s))}\right|\leqslant K_{2}.$
In addition, it follows from Lemma 2.3 that for $s\in L^{\pm}$ and $n\in\mathbb{Z}_{\geqslant 0}$,
$\prod_{j=1}^{n}\left|\frac{-d+j-s}{-b+j-s}\right|=\mathcal{O}\left(\left(n+1\right)^{2|b-d|}\right).$
In view of the fact that
$h_{n}(s)=h_{0}(s)\prod_{j=1}^{n}\frac{d-j+s}{b-j+s}$
and the estimates above, we get
$\left|R_{\pm}(z)\right|$ $\leqslant C\left(n+1\right)^{2|b-d|}\left(\int_{s\in L^{\pm},|s|<K_{1}}|h_{0}(s)z^{s}||ds|+\int_{s\in L^{\pm},|s|\geqslant K_{1}}|h_{0}(s)z^{s}||ds|\right)$
$\leqslant C\left(n+1\right)^{2|b-d|}\left(K_{3}+K_{2}^{2}\int_{s\in L^{\pm},|s|\geqslant K_{1}}\left|\Gamma(-s+\alpha)z^{s}\right||ds|\right).$
According to [11, Lemma 2.8], the contours $L^{\pm}$ can be deformed to pass through the saddle point $s=-z$ and to be bounded away from the points shown in (2.4). Hence, the integral in the last line has the order of magnitude $\mathcal{O}\big{(}|z|^{\Re(\alpha)}\mathrm{e}^{-\Re(z)}\big{)}$.
The expansion now (2.2) follows from the estimates of $R_{v}(z)$ and $R_{\pm}(z)$. ∎
The second provides expansions of ${}_{2}F_{2}[a,b+n;c,d+n;z]$ for large $z\in\mathbb{C}\setminus\mathbb{R}_{\geqslant 0}$.
Theorem 2.5.
Assume that $a,b\in\mathbb{C}$ and $c,d\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Then for $n\in\mathbb{Z}_{\geqslant\max\{0,-\Re(a),-\Re(b)\}}$,
${}_{2}F_{2}\left[\begin{matrix}a,b+n\\ c,d+n\end{matrix};-z\right]=\left(n+1\right)^{\max\{0,\Re(d-b)\}}\mathcal{O}(1)+\left(n+1\right)^{2|b-d|}\mathcal{O}\left(\left|z\right|^{\Re(a+b-c-d)}\mathrm{e}^{-\Re(z)}\right)$
as $z\to\infty$ such that $\left|\arg(z)\right|<\pi$ and $z$ is bounded away from the points $-b+k\,(k\in\mathbb{Z})$.
Proof..
It is sufficient to observe that
${}_{2}F_{2}\left[\begin{matrix}a,b+n\\ c,d+n\end{matrix};-z\right]=\frac{\Gamma(c)\Gamma(d+n)}{\Gamma(a)\Gamma(b+n)}R_{-n,0}(z),$
where $R_{-n,0}(z)$ is defined in (2.3) (The contour is suitably indented in order to avoid the pole $s=0$ of $\Gamma(-s)$). The result follows by repeating the proof of Theorem 2.4. ∎
The third establishes explicit upper bounds of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for $z\to+\infty$ and refines the rough bounds in [4, Lemma 3.3].
Theorem 2.6.
Let $p:=\Re(a-c)+\max\{0,\Re(b-d)\}$.
(i) Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $d\in\mathbb{C}\setminus\mathbb{Z}$. Then for $n\in\mathbb{Z}_{\geqslant 0}$,
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};z\right]=\mathcal{O}\left(\left(n+1\right)^{2|b-d|}z^{p}\mathrm{e}^{z}\right),\quad z\to+\infty.$
(ii) Assume that $a,b\in\mathbb{C}$ and $c,d\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Then for $n\in\mathbb{Z}_{\geqslant 0}$,
${}_{2}F_{2}\left[\begin{matrix}a,b+n\\ c,d+n\end{matrix};z\right]=\mathcal{O}\left(\left(n+1\right)^{2|b-d|}z^{p}\mathrm{e}^{z}\right),\quad z\to+\infty.$
Proof..
By following the proof of [4, Lemma 3.3], the results follow from Lemma 2.3. ∎
3.  Proof of the main results
We are now in a position to prove our main results.
Proof of Theorem 1.2.
Recall the series representation (1.4). We have $V_{1}(x,y)=\mathcal{O}(1)$ since $\frac{y}{1-x}$ is bounded. Hence the main contribution of $\Psi_{1}$ comes from $V_{2}(x,y)$. Now divide $V_{2}(x,y)$ into two parts:
$V_{2}(x,y)=\sum_{n=0}^{N-1}+\sum_{n=N}^{\infty}=:V_{2}^{(L)}(x,y)+V_{2}^{(R)}(x,y),$ (3.1)
where $N$ is a positive integer to be determined.
For $10\leqslant n\leqslant N-$ 1, recall the exponential expansion of ${}_{2}F_{2}$ [8, Equation (5.8)]
${}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]=\frac{\Gamma(c)\Gamma(d-n)}{\Gamma(a)\Gamma(b-n)}\mathrm{e}^{y}\sum_{k=0}^{N-1}c_{k,n}y^{a+b-c-d-k}+\mathcal{O}\left(y^{\Re(a+b-c-d)-N}\mathrm{e}^{y}\right),$ (3.2)
as $y\to+\infty$, where the coefficients $c_{k,n}$ are given by [12, Equation (12)]
$c_{k,n}:=\frac{\left(c+d-a-b\right)_{k}\left(n+1-b\right)_{k}}{k!}\,_{3}F_{2}\left[\begin{matrix}-k,c-a,d-a-n\\ c+d-a-b,b-n-k\end{matrix};1\right].$ (3.3)
Inserting (3.2) into (3.1), we can infer that
$V_{2}^{(L)}(x,y)=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{n=0}^{N-1}\frac{\left(b\right)_{n}}{n!}\left(1-x\right)^{-n}\sum_{k=0}^{N-1}c_{k,n}^{*}y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-N}\mathrm{e}^{y}\right),$
where
$c_{k,n}^{*}=\frac{\left(b-a+c^{\prime}\right)_{k}\left(n+b-a+1\right)_{k}}{k!}\,_{3}F_{2}\left[\begin{matrix}-k,-n,c+c^{\prime}-a-1\\ b-a+c^{\prime},a-b-n-k\end{matrix};1\right].$
Hence
$V_{2}^{(L)}(x,y)$ $=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}y^{-k}\sum_{j=0}^{k}c_{k-j,j}^{*}\frac{\left(b\right)_{j}}{j!}\Big{(}\frac{y}{1-x}\Big{)}^{j}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)\left(\frac{1-x}{y}\right)^{b}y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-N}\mathrm{e}^{y}\right),$ (3.4)
where $a_{k}(x,y)$ is given by (1.5).
For $n\geqslant N$, a combination of (2.1) and Theorem 2.6 yields
$\left|V_{2}^{(R)}(x,y)\right|$ $\leqslant Cy^{p}\mathrm{e}^{y}\sum_{n=N}^{\infty}\left(n+1\right)^{q}\left|1-x\right|^{-n}\leqslant Cy^{p}\mathrm{e}^{y}\sum_{n=N}^{\infty}\left(n+1\right)^{q}\gamma_{2}^{n}y^{-n}$
$=C\gamma_{2}^{N}y^{p-N}\mathrm{e}^{y}\sum_{k=0}^{\infty}\left(k+N+1\right)^{q}\gamma_{2}^{k}y^{-k}=\mathcal{O}\left(y^{p-N}\mathrm{e}^{y}\right),$ (3.5)
where $p=\Re(a-c-c^{\prime}+1)+\max\{0,\Re(c-b-1)\}$ and $q=\Re(c)-2+2|b-c+1|$.
For any given positive integer $M$, take $N=M+\lceil p-\Re(a-b-c^{\prime})\rceil$. Since $p\geqslant\Re(a-b-c^{\prime})$, one has $N\geqslant M$ and $\Re(a-b-c^{\prime})-M\geqslant p-N$. Then it follows from (3.4) and (3.5) that
$\left(1-x\right)^{-b}V_{2}(x,y)$ $=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-2b-c^{\prime})-N}\mathrm{e}^{y}\right)+\mathcal{O}\left(y^{p-\Re(b)-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-2b-c^{\prime})-M}\mathrm{e}^{y}\right)+\mathcal{O}\left(y^{p-\Re(b)-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-2b-c^{\prime})-M}\mathrm{e}^{y}\right),$
which completes the proof. ∎
Proof of Theorem 1.3.
Estimates of error terms here are similar to those of Theorem 1.2, so we omit the details and just extract the expansions.
Let us start with a convergent series
$F(x,y):=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}}\,_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]\left(1-x\right)^{-n}.$
Recall the asymptotic expansion of ${}_{2}F_{2}(y)$ [8, Equation (5.8)]: for $0\leqslant n\leqslant N$,
$\begin{split}{}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]={}&\frac{\Gamma(c)\Gamma(d)}{\Gamma(a)\Gamma(b)}\frac{\left(a_{3}\right)_{n}}{\left(b_{3}\right)_{n}}\bigg{\{}S_{n}(-y)+T_{n}(-y)+\mathrm{e}^{y}\sum_{k=0}^{N-1}c_{k,n}y^{a+b-c-d-k}\\ &+\mathcal{O}\left(\left|y\right|^{-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a+b-c-d)-N}\mathrm{e}^{\Re(y)}\right)\bigg{\}},\end{split}$ (3.6)
where $a_{3}=1-b,\,b_{3}=1-d$, $S_{n}(-y)$ and $T_{n}(-y)$ are defined in Theorem 2.4, and the coefficients $c_{k,n}$ are given by (3.3). For $n\geqslant N+1$, recall the expansion (2.2). Using (2.1) and the definition of $F(x,y)$ then gives
$F(x,y)=\frac{\Gamma(c)\Gamma(d)}{\Gamma(a)\Gamma(b)}\left(S_{w}(x,y)+T_{w}(x,y)+E_{N}(x,y)\right)+\mathcal{O}\left(\left|y\right|^{-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a+b-c-d)-N}\mathrm{e}^{\Re(y)}\right),$ (3.7)
where
$S_{w}(x,y)$ $:=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}}S_{n}(-y)\left(1-x\right)^{-n},$
$T_{w}(x,y)$ $:=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}}T_{n}(-y)\left(1-x\right)^{-n},$
and
$E_{N}(x,y):=y^{a+b-c-d}\mathrm{e}^{y}\sum_{k=0}^{N-1}y^{-k}\sum_{j=0}^{k}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}c_{k-j,j}\left(\frac{y}{1-x}\right)^{j}.$
Direct computation gives
$S_{w}(x,y)={}$ $\frac{\Gamma(a)\Gamma(b-a)}{\Gamma(c-a)\Gamma(d-a)}\sum_{k=0}^{\lfloor w-\Re(a)\rfloor}\frac{\left(a\right)_{k}\left(a-c+1\right)_{k}\left(a-d+1\right)_{k}}{\left(1\right)_{k}\left(a-b+1\right)_{k}}$
$\qquad\cdot\left(-y\right)^{-a-k}\,_{5}F_{4}\left[\begin{matrix}a_{1},a_{2},a_{3},1,a-d+k+1\\ b_{1},b_{2},b_{3},a-b+k+1\end{matrix};\frac{1}{1-x}\right]$
[…]
where
$s_{k}(x,y)=\sum_{j=0}^{k}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}\frac{\left(a\right)_{k-j}\left(a-c+1\right)_{k-j}}{\left(1\right)_{k-j}}\left(\frac{y}{x-1}\right)^{j}.$
Write $K=\lfloor w-\Re(b)\rfloor$ and $\gamma=\frac{y}{1-x}$. Then
$T_{w}(x,y)$ $=\left(-y\right)^{-b}\sum_{n=0}^{\infty}\sum_{k=0}^{n+K}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}}\frac{\Gamma(b-n+k)\Gamma(a-b+n-k)}{\Gamma(d-b-k)\Gamma(c-b+n-k)}\frac{\left(-1\right)^{k}}{k!}\left(-y\right)^{n-k}\left(1-x\right)^{-n}$
$=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\left(-y\right)^{-b}\sum_{n=0}^{\infty}\sum_{k=0}^{n+K}t_{n,k}\gamma^{n}\left(-y\right)^{-k},$
where
$t_{n,k}:=\frac{\left(b\right)_{k}\left(b-c+1\right)_{k}\left(b-d+1\right)_{k}}{\left(b-a+1\right)_{k}k!}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}\left(a-b-k\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}\left(1-b-k\right)_{n}\left(c-b-k\right)_{n}}.$
Now split $T_{w}(x,y)$ into two parts:
$T_{w}(x,y)=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\left(-y\right)^{-b}\left(T_{w}^{-}(x,y)+T_{w}^{+}(x,y)\right),$
where
$T_{w}^{-}(x,y)$ $=\sum_{n=0}^{\infty}\sum_{k=0}^{n}t_{n,k}\gamma^{n}\left(-y\right)^{-k}=\sum_{k=0}^{\infty}\sum_{n=k}^{\infty}t_{n,k}\gamma^{n}\left(-y\right)^{-k}$
$=\sum_{k=0}^{K}\sum_{n=k}^{\infty}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-K-1}\right)$
and
$T_{w}^{-}(x,y)$ $=\sum_{n=0}^{\infty}\sum_{k=n+1}^{n+K}t_{n,k}\gamma^{n}\left(-y\right)^{-k}$
$=\sum_{n=0}^{K}\sum_{k=n+1}^{n+K}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-K-1}\right)$
$=\sum_{k=1}^{K}\sum_{n=0}^{k-1}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-K-1}\right).$
Since the inner sum in the last line is null when $k=0$, we derive
$T_{w}(x,y)$ $=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\left(-y\right)^{-b}\sum_{k=0}^{K}\sum_{n=0}^{\infty}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-\Re(b)-K-1}\right)$
$=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\sum_{k=0}^{K}t_{k}(x,y)\left(-y\right)^{-b-k}+\mathcal{O}\left(\left|y\right|^{-\Re(b)-K-1}\right),$
where
$t_{k}(x,y)=\frac{\left(b\right)_{k}\left(b-c+1\right)_{k}\left(b-d+1\right)_{k}}{\left(b-a+1\right)_{k}k!}\,_{5}F_{5}\left[\begin{matrix}a_{1},a_{2},a_{3},1,a-b-k\\ b_{1},b_{2},b_{3},1-b-k,c-b-k\end{matrix};\frac{y}{1-x}\right].$
The asymptotic expansion of $\Psi_{1}$ follows from (1.4) and the expansion (3.7). ∎
4.  Concluding remarks
We have established the complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments. Our derivation is based on the estimates of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for large $z$ (see Theorems 2.42.6), but the exponential expansions are not included. We conjecture that the more accurate expansion (3.6) is valid for $n\in\mathbb{Z}$ with remainders multiplied by $n^{\lambda}$ for some $\lambda>0$. Braaksma’s work [1] seems helpful to examine this conjecture.
Regarding asymptotics of $\Psi_{1}$ for large $y$, it was partially studied in [5, Section 3.1]. The series representations (1.4) and [4, Equation (3.4)] appear to provide complete asymptotic expansions for large $y$ with unrestricted $\arg(x)$ and $\arg(y)$. We will determine appropriate methods to verify this. Additionally, the asymptotics of other Humbert functions will be explored in future studies.
Here we would like to make further remarks on our Theorems 1.2 and 1.3. Note that $\Psi_{1}$ has a Kummer-type transformation [3, Equation (2.54)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\left(1-x\right)^{-a}\Psi_{1}\left[a,c-b;c,c^{\prime};\frac{x}{x-1},\frac{y}{1-x}\right],$ (4.1)
which elucidates that the singularity of $\Psi_{1}[x,y]$ as $x\to 1$ is equivalent to its asymptotics under the condition (1.6). As an example, one may get from (4.1) and Theorem 1.2 that for fixed $y>0$,
$\Psi_{1}\left[1,\frac{1}{2};\frac{3}{2},\frac{1}{2};x,y\right]\sim\frac{\sqrt{\pi}}{2}y^{-\frac{1}{2}}\left(1-x\right)^{\frac{1}{2}}\mathrm{e}^{\frac{y}{1-x}},\quad x\to 1^{-},$ (4.2)
which is confirmed by [4, Equation (3.4)] and was mentioned in [6, p. 22].
Finally, by using Mathematica 12, we provide a numerical verification of Theorems 1.2 and 1.3. The value of $\Psi_{1}=\Psi_{1}\bigl{[}1,\frac{1}{2};\frac{1}{3},\frac{1}{4};tx,ty\bigr{]}$ for large $t>0$ is evaluated by using (1.2), and the corresponding leading term is denoted by $L_{\Psi_{1}}$. Table 1 illustrates that the ratio $\frac{\Psi_{1}}{L_{\Psi_{1}}}$ approaches unity as $t\to+\infty$.
Table 1: Numerical verification of $\Psi_{1}\bigl{[}1,\frac{1}{2};\frac{1}{3},\frac{1}{4};tx,ty\bigr{]}$.
$x$ $y$ $t$ $\frac{\Psi_{1}}{L_{\Psi_{1}}}$
$-1$ $2$ $10$ $0.971796$
$-1$ $2$ $100$ $0.997355$
$-1$ $2$ $1000$ $0.999737$
$-1$ $-2$ $10$ $1.045341$
$-1$ $-2$ $100$ $1.004387$
$-1$ $-2$ $1000$ $1.000438$
Acknowledgements
We would like to thank Malte Henkel for suggesting the inclusion of the example (4.2), and Nico Temme for bringing Braaksma’s work to our attention in a personal discussion with the first author.